Suppose that 12% of adults are classified as hypertensive. For people that are truly hypertensive the probability of a positive test is 0.85. Also, for people that are not truly hypertensive the probability of a positive test is 0.24. What is the probability that an individual is not hypertensive?

A test on probability is to be given next Friday. Suppose that 80 percent of the students study for the test and 20 percent do not. If a student studies for the exam, the probability that he or she will pass is 0.75. If the student does not study, the probability that he or she will pass is 0.20. What is the probability he or she really studied and passes the test?

Given five women and three men, in how many ways can a three-member committee be selected (blank) with no restriction, (blank) should be exactly one woman in the committee. Fill the blanks.
Answer #1

Ans:

1)

P(hypertensive)=0.12

P(positive/hypertensive)=0.85

P(positive/not hypertensive)=0.24

P(not hypertensive)=1-0.12=0.88

P(positive)=0.85*0.12+0.24*0.88=0.3132

P(hypertensive/positive)=P(positive/hypertensive)*P(hypertensive)/P(positive)

=0.85*0.12/0.3132=0.3257

2)

P(pass/study)=0.75

P(pass/not study)=0.20

P(study)=0.8

P(not study)=0.2

P(study and pass)=P(pass/study)*P(study)=0.75*0.80=0.6